At a concession stand, one person purchased 3 drinks and 3 pretzels for $9.00.
Another person buys 4 drinks and 2 pretzels for $8.50. What is the price of one drink?

Answer by lwsshak3(11628)

At a concession stand, one person purchased 3 drinks and 3 pretzels for $9.00.
Another person buys 4 drinks and 2 pretzels for $8.50. What is the price of one drink?
let d=number of drinks
let p=number of pretzels
3d+3p=9.00
4d+2p=8.50
6d+6p=18.00 (mult. 1st eq by 2
12d+6p=25.50 (mult. 2nd eq by 3
subtract
-6d=-7.50
d=1.25
3p=9.00-3d=9-3.75
p=1.75
price of one drink=$1.75